OpenStudy (anonymous):

if f(x)=1-x^2, what is the range of f?

OpenStudy (anonymous):

First wat is the Domain can u find out

OpenStudy (anonymous):

i dont know the domain

OpenStudy (anonymous):

y = 1 - x^2 x^2 = 1-y x = sqrt 1-y

OpenStudy (anonymous):

so it should not be a complex number

OpenStudy (anonymous):

the final answer is f(x) is greater than or equal to 1

OpenStudy (anonymous):

1 - y > =0 y < = 1

OpenStudy (anonymous):

Range = [1 , infinity)

OpenStudy (anonymous):

Did u understand

OpenStudy (anonymous):

no how do you get that answer?

OpenStudy (anonymous):

take y = f(x)

OpenStudy (anonymous):

okay then what?

OpenStudy (anonymous):

f(x)=1-x^2

OpenStudy (anonymous):

y = 1 - x^2 x^2 = 1-y x = sqrt 1-y

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

ok.....

OpenStudy (anonymous):

u have more question on this topic??

OpenStudy (anonymous):

just inform me